if 2000 is 14%, what is the other 86%. I'm just old and forgot how to do this?
Found 2 solutions by dragonwalker, richwmiller:
Answer by dragonwalker(73)

Well, think about it this way. 14% represents 14 units of 100 (i.e. percent) so you want to find out what 1% is.
To do this divide 2000 by 14 = 2000/14
= 142.86
so if 1 % is 142.86 then 86 % is 86 x 142.86
= 12285.96
= 12286
You could multiply 142.86 by 100 to get the total amount: = 14286
this makes sense as 2000 + 12286 gives the same total.

Answer by richwmiller(17219)

.14x=2000
x=2000/.14
x=200000/14
x=100000/7=14285.71428571429
14285.71428571429-2000=12285.71428571429= 86% x